We work at the Carlton Carpet Company. The average number of defects on a 10’ x 10’ piece of carpet is 5. What is the probability of at least 2 defects?


1
Expert's answer
2021-08-05T16:25:57-0400

This is a Poisson's distribution with "\lambda=5."

"P(X\ge2)=1-P(X=0)-P(X=1)="

"=1-e^{-5}\frac{5^0}{0!}-e^{-5}\frac{5^1}{1!}=1-6e^{-5}=0.9596."